OpenStudy (anonymous):

i write the wrong question 2(x-2)^2-6 find allints in exact form

OpenStudy (anonymous):

very similar to the last one

OpenStudy (anonymous):

replace x by 0 to get the y intercept set y = 0 and solve for x to get the x intercepts

OpenStudy (anonymous):

y=2

OpenStudy (anonymous):

but i need help with the x's

OpenStudy (anonymous):

yes y = 2

OpenStudy (anonymous):

\[2(x-2)^2-6=0\] \[2(x-2)^2=6\] \[(x-2)^2=3\] \[x-2=\pm\sqrt{3}\] \[x=2\pm\sqrt{3}\]

OpenStudy (anonymous):

make sure you follow all the steps. they are all there i skipped nothing

OpenStudy (anonymous):

so you dont have to expand the brackets??

OpenStudy (anonymous):

oh heck no

OpenStudy (anonymous):

it is just in the form i want it. check the steps. if you expand you have way more work to do after you expand, so you would be working against yourself

OpenStudy (anonymous):

yeh cos i was expanding and going WTF this clears it up just 1 question why is it \[\pm\] not just +

OpenStudy (anonymous):

because if you have \[x^2=25\] you have two choices for x \[x=5\] or \[x=-5\] so it is an abbreviation for the two choices

OpenStudy (anonymous):

ohn k thanks'

OpenStudy (anonymous):

\[(x-2)^2=25\] \[x-2=5\] or \[x-2=-5\] \[x=7\] or \[x=-3\]

OpenStudy (anonymous):

\[(x-2)^2=3\] \[x-2=\sqrt{3}\] or \[x-2=-\sqrt{3}\] \[x=2+\sqrt{3}\] or \[x=2-\sqrt{3}\]